# Protect Before Generate: Error Correcting Codes within Discrete Deep Generative Models

### v1

@@ Appendix I Evaluating log-likelihood using the soft-encoding model
The reparameterization trick introduced in (5) requires that bits are independent, a condition not met in $\boldsymbol{c}$ once the code’s structure is introduced. To address this issue, during training, we employ a soft encoding strategy. We assume the bits in $\boldsymbol{c}$ are independent and equally distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. Therefore, instead of directly repeating sampled bits in $\boldsymbol{m}$ to obtain $\boldsymbol{c}$ following $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, we repeat the posterior probabilities for the copies of the same bit and sample $\boldsymbol{z}$ using the reparameterization trick in (5). Remark that, despite the soft encoding assumption during training, the generative results in Fig. 4, 6, 13, 19, 24, 24, and 29 are obtained through hard encoding. Namely, we sampled an information word $\boldsymbol{m}$ and obtained $\boldsymbol{c}$ by repeating its bits.
@@
While the hard-encoding images are visually appealing, to evaluate the coded DVAE LL for a given image $\boldsymbol{x}$, we have to leverage the soft-encoding model since it is unrealistic for a sample from the soft-encoding model to produce equally repeated bits. In the soft-coded model, we sample bit probabilities from a prior distribution that we model through the product of $M$ independent Beta distributions, and we use a proposal distribution model similar to the Vamp-prior in [28]. Namely, a mixture model with components given by $q(\boldsymbol{m}|\boldsymbol{x})$ for different training points.
@@
In the main text and Section H, we report the log-likelihood values for different model configurations trained on the FMNIST dataset. Table 4 presents the results obtained for the MNIST dataset. Due to their simplicity, these datasets do not require high-dimensional latent spaces, and competitive results can be achieved with just 8 or 10 information bits. However, for more complex datasets like CIFAR10 or Tiny ImageNet, high-dimensional latent spaces are necessary to capture spatial information, colors, and textures. For these high-dimensional datasets, we were unable to obtain valid log-likelihood estimates because the number of samples needed for importance sampling to converge was too large. However, given the difference in the level of detail in the reconstructed and generated images between coded and uncoded models (see Sections D and E), coded models are expected to better approximate the true marginal likelihood of the data.
Appendix J Variational inference at codeword level
@@
The KL Divergence term is approximated via Monte Carlo, drawing samples from $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$. As it is not possible to backpropagate through discrete variables, we approximate the gradients with respect to the parameters of the encoder using the REINFORCE leave-one-out estimator [23, 13], given by
@@
Defining a distribution over the codebook can seem intuitive, but scalability becomes challenging as the size of the codebook increases. The reason is that the posterior distribution must be evaluated for all codewords during both inference and test time. However, it can still provide a bound, enabling the utilization of more complex codes with theoretical guarantees that can outperform the previously proposed repetition codes.
@@
Appendix K Computational resources
@@
Appendix L Coded DVAE scheme
Figure 39: Graphic representation of the coded DVAE, illustrating both the generative and inference paths.
Appendix M Hierarchical coded DVAE scheme
Figure 40: Graphic representation of the hierarchical coded DVAE, illustrating both the generative and inference paths.

### v2

@@ Appendix I Beyond Repetition Codes
The reparameterization trick introduced in (5) requires that bits are independent, a condition not met in $\boldsymbol{c}$ once the code’s structure is introduced. To address this issue, during training, we employ a soft encoding strategy. We assume the bits in $\boldsymbol{c}$ are independent and equally distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. Therefore, instead of directly repeating sampled bits in $\boldsymbol{m}$ to obtain $\boldsymbol{c}$ following $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, we repeat the posterior probabilities for the copies of the same bit and sample $\boldsymbol{z}$ using the reparameterization trick in (5). Remark that, despite the soft encoding assumption during training, the generative results in Fig. 4, 8, 15, 21, 26, 26, and 31 are obtained through hard encoding. Namely, we sampled an information word $\boldsymbol{m}$ and obtained $\boldsymbol{c}$ by repeating its bits.
@@
While the hard-encoding images are visually appealing, to evaluate the Coded-DVAE LL for a given image $\boldsymbol{x}$, we have to leverage the soft-encoding model since it is unrealistic for a sample from the soft-encoding model to have equally repeated bits. In the soft-coded model, we sample bit probabilities from a prior distribution that we model through the product of $M$ independent Beta distributions, and we use a proposal distribution model similar to the Vamp-prior in Tomczak and Welling [2018]. Namely, a mixture model with components given by $q(\boldsymbol{m}|\boldsymbol{x})$ for different training points.
@@
In the main text and Section K, we report the LL values for different model configurations trained on the FMNIST dataset. Table 4 presents the results obtained for the MNIST dataset. Due to their simplicity, these datasets do not require high-dimensional latent spaces, and competitive results can be achieved with just 8 or 10 information bits. However, for more complex datasets like CIFAR10 or Tiny ImageNet, high-dimensional latent spaces are necessary to capture spatial information, colors, and textures. For these high-dimensional datasets, we were unable to obtain valid log-likelihood estimates because the number of samples needed for importance sampling to converge was too large. However, given the difference in the level of detail in the reconstructed and generated images between coded and uncoded models (see Sections F and G), coded models are expected to better approximate the true marginal likelihood of the data.
Appendix M Connection to previous works on VAEs as source coding methods
Effective learning of low-dimensional discrete latent representations from complex data is a technically challenging task. In this work, we propose a novel method to improve inference in discrete VAEs within a fully probabilistic framework, introducing a new perspective on the inference problem. While previous studies have analyzed VAEs using rate-distortion (RD) theory [Chen et al., 2022, Townsend et al., 2019, Van Den Oord et al., 2017], our approach stems from a different yet complementary perspective that remains compatible with these works.
In the literature, VAEs are often interpreted as lossy compression models (e.g., VAEs as source coding methods), but our contribution is best understood from a generative perspective. We conceptualize inference via $q(\boldsymbol{m}|\boldsymbol{x})$ as a decoding process, where the goal is to recover the discrete latent variable from the observed data. We sample a latent vector $\boldsymbol{m}$, generate a data point $\boldsymbol{x}$ (e.g., an image), and aim to minimize the error rate in recovering $\boldsymbol{m}$ from $\boldsymbol{x}$. Achieving this requires the variational approximation $q(\boldsymbol{m}|\boldsymbol{x})$ to closely match the true posterior $p(\boldsymbol{m}|\boldsymbol{x})$.We demonstrate that incorporating redundancy into the generative process reduces errors in estimating $\boldsymbol{m}$, leading to a more accurate latent posterior approximation. This improvement directly enhances our model’s overall performance, as confirmed by our experimental results. Additionally, error rates and variational gaps are related through bounds derived from mismatch hypothesis testing. These bounds demonstrate that minimizing the variational gap by maximizing the ELBO effectively tightens an upper bound on the error rate [Schlüter et al., 2013, Yang et al., 2024].
In RD theory, the primary focus is on compression within the latent space, typically analyzed from an encoder/decoder and reconstruction (distortion) perspective. RD theory establishes theoretical limits on achievable compression rates and describes how practical models may diverge from these limits. Using RD practical compression methods, Townsend et al. [2019] demonstrates how asymmetric numeral systems (ANS) can be integrated with a VAE to improve its compression rate by jointly encoding sequences of data points, bringing performance closer to RD theoretical limits. Similarly, Chen et al. [2022] shows that a complex prior distribution in a VAE using an autoregressive invertible flow narrows the gap between the approximate and the true posterior distribution, thereby enhancing the overall performance of the VAE. We note that even using an independent prior, the hierarchical code structure outlined in Section J naturally decouples information across different latent spaces at various conceptual levels. Specifically, the most relevant information (class label) is captured by the most protected latent space, while the other space captures fine-grained features. This effect is not straightforward to enforce through direct design of a more complex prior.
Our method complements all these efforts by introducing redundancy in the generative pathway to enhance variational inference, leading to more accurate latent posterior approximations. Notably, even when using a complex prior distribution, ECCs can still be leveraged to enhance inference and overall model performance. In this work, we demonstrate that our approach yields more robust models even with a simple, fixed independent prior, as evidenced by improved LL, generation quality, and reconstruction metrics. Moreover, in Section J, we show that integrating a hierarchical ECC with the same independent prior leads to even greater performance gains.
It is also important to discuss the VQ-VAE [Van Den Oord et al., 2017], one of the most relevant works related to our approach, which is well-known for effectively learning compressed discrete representations of complex data. We believe that highlighting key differences between VQ-VAEs and our method can provide useful insights. A key distinction lies in the structure and dimensionality of the latent space. In image modeling, VQ-VAEs typically employ a latent matrix where indices correspond to codewords in a codebook. This design allows different codewords to capture specific patches of the original image, improving reconstruction and generation (the latter through an autoregressive prior on the latent representations). However, representing each data point as a grid of embeddings complicates interpretability. In contrast, our method encodes the entire image into the latent space, rather than partitioning it into patches. Another major difference is the nature of the encoder. VQ-VAEs rely on a non-probabilistic encoder that maps inputs to the nearest latent code using a distance-based metric. This deterministic mapping limits the model’s ability to capture and quantify uncertainty in the latent space. Our approach, on the other hand, operates within a fully probabilistic framework, allowing for uncertainty quantification in the learned latent representations. Additionally, our method is fully differentiable, enabling seamless gradient computation and backpropagation for end-to-end training. Our approach also allows us to leverage the reparameterization trick introduced in DVAE++ [Vahdat et al., 2018b], which we found to be more stable than continuous relaxations like the Gumbel-Softmax used in VQ-VAEs with stochastic quantization [Williams et al., 2020].
Appendix N Variational inference at codeword level
@@
The KL Divergence term is approximated via Monte Carlo, drawing samples from $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$. Since it is not possible to backpropagate through discrete variables, we approximate the gradients with respect to the parameters of the encoder using the REINFORCE leave-one-out estimator [Salimans and Knowles, 2014, Kool et al., 2019], given by
@@
Defining a distribution over the codebook can seem intuitive, but scalability becomes challenging as the size of the codebook increases. The reason for this is that the posterior distribution must be evaluated for all codewords during both inference and test time. However, it can still provide a bound, enabling the utilization of more complex codes with theoretical guarantees that can outperform the previously proposed repetition codes.
@@
Appendix O Computational resources
@@
Appendix P Coded-DVAE scheme
Figure 41: Graphic representation of the Coded-DVAE, illustrating both the generative and inference paths.
Appendix Q Hierarchical Coded-DVAE scheme
Figure 42: Graphic representation of the hierarchical Coded-DVAE, illustrating both the generative and inference paths.